[figure: 9.7 cm]

THE ELECTRONIC AND ELECTROCHEMICAL PROPERTIES

OF THE LaNi_{5}-BASED ALLOYS

A. Szajek^{a}, M. Jurczyk^{b}, and M. Nowak^{b} aInstitute of Molecular Physics, Polish Academy of Sciences, Poznań Poland

bInst. of Materials Sci. and Engineering, Poznań Univ. of Technology, Poznań Poland

The battery electrode materials La(Ni − Mn − Co − Al)_{5} are substitutional derivatives of LaNi_{5} in which La occupies site 1(a), and Ni, Mn, Al, and Co sites 2(c) and 3(g) of space group P 6/mmm. Nanocrystalline LaNi_{5}-type materials have been prepared by mechanical alloying followed by annealing. The amorphous phase forms directly from the starting mixture of the elements, without other phase formation. The asmilled powders were heat treated at 1000K for 0.5h under high purity argon to form the hexagonal CaCu_{5}-type phase. The powders were examined by XRD analysis, with Co K_{α} radiation, at the various stages during milling prior to and after annealing. The electrochemical properties of these materials have been investigated and discharge capacity enhancement has been observed. The electronic structure has been studied by the tight binding version of the linear muffin-tin orbital method in the atomic sphere approximation (T B LMT O ASA). The startting LaNi_{5} is not magnetic but Al, Co, and Mn change band structure and in doped materials magnetic order is observed. The values of magnetic moments located on Mn atoms reach at leat 3μ_{B}/atom. This work was supported by the KBN Grant No. PBZ-KBN-013/T08/02

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : A. Szajek

Address for correspondence : Institute of Molecular Physics Polish Academy of Sciences ul. Smoluchowskiego 17, 60-179 Poznań, Poland

Email address : szajek@ifmpan.poznan.pl